# https://github.com/autonomousvision/sledge

### v1

@@ Abstract
SLEDGE is the first generative simulator for vehicle motion planning trained on real-world driving logs. Its core component is a learned model that is able to generate agent bounding boxes and lane graphs. The model’s outputs serve as an initial state for traffic simulation. The unique properties of the entities to be generated for SLEDGE, such as their connectivity and variable count per scene, render the naive application of most modern generative models to this task non-trivial. Therefore, together with a systematic study of existing lane graph representations, we introduce a novel raster-to-vector autoencoder (RVAE). It encodes agents and the lane graph into distinct channels in a rasterized latent map. This facilitates both lane-conditioned agent generation and combined generation of lanes and agents with a Diffusion Transformer. Using generated entities in SLEDGE enables greater control over the simulation, e.g. upsampling turns or increasing traffic density. Further, SLEDGE can support 500m long routes, a capability not found in existing data-driven simulators like nuPlan. It presents new challenges for planning algorithms, evidenced by failure rates of over 40% for PDM, the winner of the 2023 nuPlan challenge, when tested on hard routes and dense traffic generated by our model. Compared to nuPlan, SLEDGE requires 500$\times$ less storage to set up (<4GB), making it a more accessible option and helping with democratizing future research in this field.
@@ 1 Introduction
While recent breakthroughs in generative AI have revolutionized natural image synthesis [16, 4], generative models are yet to find widespread adoption in autonomous driving. In contrast to the regular pixel grid of images, self-driving planners typically require abstract bird’s eye view (BEV) representations as input which characterize the most important scene elements (e.g., lanes, traffic lights, static and dynamic objects) in a compact, vectorized format (see Fig. 2). These representations are a key component of data-driven simulators which are necessary for rigorous evaluation of planners [12, 50, 25, 11]. However, learning a generative model on such irregular vectorized representations is hard.
Consequently, many existing data-driven simulators [22, 17] are initialized by simply replaying logs of abstract representations. They extract the local lane graph from a High Definition (HD) map and object bounding boxes from pre-recorded annotation logs. They test planning algorithms by simulating short scenarios ($\sim$15 seconds long) in which the algorithms must safely traverse the route followed in the original log. However, to provide sufficient diversity among routes for comprehensive testing, these simulators require huge databases, e.g., nuPlan [22] consists of 1300 hours of driving logs which require over 2TB of storage. Such high resource requirements heighten the barrier for entry into the field of vehicle motion planning.
In this paper, we study the task of generating simulation-ready abstract representations of driving scenes. Using generative models as an alternative to log replays has the potential for significant compression [39]. Furthermore, abstract representations are far lower dimensional than visual data, simplifying the generation task. However, the unique characteristics of abstract representations in driving scenes pose new challenges to modeling them, e.g., they require accurate topological connectivity and precise modeling of geometry (e.g., parallel/perpendicular lines) in order to be useful for simulation.
@@
To tackle these complexities, we first perform a systematic study of the existing representations of lane graphs used in autonomous driving. We then propose a novel representation based on a raster-to-vector autoencoder (RVAE). It represents a driving scene with a fixed-size BEV rasterized latent map (RLM). We learn to generate these RLMs with a Diffusion Transformer (DiT) [35]. Our model generates high-fidelity results enabling both lane-conditioned agent generation or joint lane and agent generation within a single flexible and scalable framework (Fig. 1). Independent to its use of generative models, SLEDGE also provides the previously missing functionality of simulating only agents within a certain radius of the ego-vehicle. By doing so, we can test on routes that are significantly longer than those currently used for evaluating planners. We find previously ignored failure modes of the state-of-the-art PDM-Closed planner [12], which is unable to complete 25-50% of our new 500m long test routes despite having failure rates below 10% on existing benchmarks.
Contributions. (1) We formalize the task of abstract scene generation for autonomous driving with a challenging benchmark and corresponding metrics. (2) We perform a systematic exploration of modern generative models (with various architectures and representations) and propose a novel latent diffusion model for synthesizing abstract driving scenes that largely outperforms other baselines. (3) We present a simulation framework, SLEDGE, which is nearly 3 orders of magnitude more storage efficient than nuPlan yet enables more rigorous testing of motion planning algorithms. SLEDGE will be made publicly available.
@@ 2 Related Work
Diffusion Models. Best known for their success in generative modeling of images [38, 16] and video [3, 4, 48], diffusion models have recently found widespread adoption in diverse domains, including point clouds [30, 49, 33], floor plans [41, 7], molecules [47], robot policies [9], traffic patterns [51], and many others. Scenario Diffusion [36], a pioneering approach in generating vehicles conditioned on HD maps with diffusion, is the closest existing approach to ours. This method uses latent diffusion with a raster-based vehicle decoder unlike our proposed transformer decoder head. Importantly, we offer significantly increased capabilities in comparison: generation of lane graphs (enabling settings both with and without map conditioning), support for pedestrians, static obstacles, and traffic lights, as well as long-horizon simulation environments with reactive agents.
@@
Generating Lane Graphs. Lane graphs, being the most important component of HD maps, are well-studied in the autonomous driving domain. They are often constructed through an expensive offline mapping process often involving human annotators [15]. However, a surge of recent work on predicting lane graphs from sensor data [26, 24] has sparked interest on generative modeling of these lane graphs. The first and only existing research study on this task, HDMapGen [32], proposes an autoregressive approach for generating lane graphs node-by-node [10]. Our experiments show that this achieves reasonable results, but is unable to match the high quality and scalability offered by diffusion models. Unlike HDMapGen, our approach jointly generates agents with lanes, and efficiently generates all scene elements in parallel. A concurrent project, DriveSceneGen [42], generates lanes and vehicles with image-space diffusion. Our approach covers agent types beyond vehicles which are critical for simulation, uses less heuristics, and is more efficient via the use of compact latent representations. Additionally, we show the successful integration of our model into a simulator, which we then use for the downstream task of evaluating planning algorithms.
@@
Data-driven Simulation. Developing an autonomous driving system necessitates rigorous testing which is costly and risky if conducted in the real world. Driving simulators are a valuable alternative [46, 14, 22]. However, simulators face challenges in ensuring realism while initializing traffic scenes, simulating traffic, and providing sensor data, for which they primarily depend on rule-based solutions [8]. Instead, data-driven simulators address these challenges by replaying traffic scenes from real-world recordings [1, 22, 23, 17]. The simulator can mine specific situations or even optimize the initial parameters for safety-critical scenarios [13, 18, 45]. Leveraging modern generative models, we take a step further and learn the underlying distribution of the real-world data. This leads to a flexible, compact and controllable simulator that goes beyond the replay-based nuPlan framework [22, 5] which we build upon.
3 SLEDGE
@@
Our goal is to design a driving scene synthesis framework that can be trained using real-world driving logs and incorporated into SLEDGE, our generative simulator. We base this framework on LDMs [38] as: (1) latent diffusion shows excellent training stability and scalability with compute. (2) One can easily construct a fixed-size latent space for diffusion that can be mapped to the variable sized set representation for simulation (Section 3.1) using detection-based transformer architectures [6, 26]. Our LDM is trained in two stages: an autoencoder (Section 3.2) followed by a diffusion model (Section 3.3). We detail the simulation of scenes generated by the LDM in Section 3.4.
@@
We represent a scene state abstractly in the format native to the nuPlan simulator. The representation combines several sets of entities, as outlined below.
@@ 3.2 Raster-to-Vector Autoencoder
Each entity type in $\mathcal{S}$ is unique. To maintain overall scene consistency, we would like to model them with a single architecture, instead of creating several independent entity-specific generative models. Furthermore, most existing tools in the literature have been developed and optimized for 2D input domains [38]. To this end, we propose the raster-to-vector autoencoder (RVAE) which unifies all entity types in $\mathcal{S}$ into a compact, shared 2D representation well-suited for diffusion modeling. The RVAE is inspired by work in object detection [6] and online mapping for autonomous driving [26].
@@
Figure 3: Raster-to-Vector Autoencoder (RVAE). We represent scenes with a rasterized latent map (RLM) consisting of two channel groups. The ‘Lanes’ group is decoded into lane segments and the ‘Agents’ group into all other scene entities, via a transformer decoder with attention masking. The RVAE is trained to predict polylines, bounding boxes, and the ego velocity in a simulation-compatible vectorized format.
Vectorization. This step is necessary to decode the unified RSI representation $\mathbf{I}$ back into the per-entity attributes (e.g., polylines, bounding boxes). We use a learned vectorization pipeline consisting of a raster encoder $\pi$ and vector decoder head $\phi$. The ResNet-50 [19] encoder takes the RSI and outputs a rasterized latent map (RLM) $\mathbf{M}=\pi(\mathbf{I})$ of shape $W^{\prime}\times H^{\prime}\times C$, where $H^{\prime}=H/2^{d}$ and $W^{\prime}=W/2^{d}$ for a downsampling factor $d\in\mathbb{N}$. $C$ is a chosen channel dimension. In practice, the RLMs we use are compact, $H^{\prime}=W^{\prime}=8$ and $C=64$. Additionally, as shown in Fig. 3, we split the RLM’s channels into 2 groups, $C=C_{L}+C_{A}$ for the lanes and agents respectively. For each group, we tokenize the latent vectors spatially, resulting in $W^{\prime}\times H^{\prime}$ lane tokens and $W^{\prime}\times H^{\prime}$ agent tokens. Following the DETR [6] paradigm, we use these tokens as keys and values for our transformer decoder $\phi$. The decoder uses a fixed number of learnable queries of each entity type, which we cap to a maximum count per entity, based on statistics from our dataset. The final decoder layer is unique per entity type and outputs the attributes specific to that entity, e.g. a $20\times 2$ set of point coordinates for a polyline, or a 6-dimensional descriptor (2D position, orientation, 2D extent, and speed) for a bounding box. It also outputs an existence attribute $p\in[0,1]$ for both polylines ($p_{line}$) and bounding boxes ($p_{box}$), which is used to handle variable counts of ground truth entities with a fixed number of queries.
@@
Training. The RVAE is optimized using both reconstruction and existence losses on the decoder outputs, and a KL divergence loss on the RLM. For reconstruction, we first match generated and ground truth entities using the Hungarian algorithm, as in [6]. We use a matching score based on the L1 error of the entity’s position attributes. We then use the L1 error summed over all attributes and averaged over all matches as the training loss. For the existence variable, we use a binary cross entropy loss based on whether the query was matched to a ground truth entity. We provide additional details in the supplementary material.
@@ 3.3 Diffusion Transformer
Training. For each scene, we sample a noise scaling factor $\sigma$ from a log-normal distribution and create a noisy sample $\hat{\mathbf{M}}=\mathbf{M}+\sigma\boldsymbol{\mathcal{E}}$, with $\boldsymbol{\mathcal{E}}$ being sampled from a standard normal distribution of the same shape as $\mathbf{M}$. We model $\delta(\hat{\mathbf{M}};\mathbf{c},\sigma)$ as a Transformer [44], following DiT [35], where $\mathbf{c}$ is a conditioning vector. For our experiments, we choose $\mathbf{c}$ to be a one-hot class label indicating the city to which the example belongs, which is readily available in nuPlan. However, it is easily possible to adapt our framework to other conditioning, e.g., images, text descriptions, or learned cluster labels based on RVAE features. The DiT architecture is simple, scalable, and free from down- or upsampling operations, making it compatible with RLMs of any spatial resolution. It applies a series of self-attention blocks to the tokenized input $\hat{\mathbf{M}}$, with the conditioning on $\mathbf{c}$ and $\sigma$ implemented using AdaLN-Zero [35]. To train the model, we optimize the L2 reconstruction loss between $\boldsymbol{\mathcal{E}}$ and $\delta(\hat{\mathbf{M}};\mathbf{c},\sigma)$.
@@ 4.1 Lane Graph Representations
Baselines. We consider three baseline representations. (1) RSI: we use the skan library [34] to extract vector polylines from the $256\times 256\times 2$ RSI representation of the lane graph. Skan is a highly optimized image processing pipeline for graph extraction (details and visualizations in supplementary). (2) RLM w/o mask: we train the RVAE without the channel group masking proposed in Section 3.2, which entangles information about agents and lanes into a single $8\times 8\times 64$ latent map instead of an $8\times 8\times 32$ tensor for the lane graph and an independent $8\times 8\times 32$ tensor for agents. (3) Vector: As an upper bound, we additionally compute our metrics for the respresentation used as target labels for the RVAE during training. This is a set of polylines shaped $N\times 20\times 2$, where we cap $N$ at 30 in our experiments. Note that the ground truth for evaluation in this experiment uses all polylines in the scene (which is sometimes greater than 30).
@@
Table 2: 64m×$\times$64m Lane Graph Generation. We show the route lengths and various Frechet distances between samples from the validation set and model outputs. We additionally include qualititve results (more examples in supplementary material). Latent diffusion combining DiT with an RLM obtains the best results. *Trained with $\sim$6$\times$ more compute than others, which already converge at the lower compute budget.
Arch. Repr. Route Frechet (CNN) $\downarrow$ Frechet (Urban Planning) $\downarrow$
Length $\uparrow$ ImNet RVEnc Conne. Densi. Reach Conve.
VAE RSI 2.68 $\pm$ 3.66 21.06 14.77 9.45 0.99 2.86 13.06
RVAE Vector 23.79 $\pm$ 9.96 9.27 14.52 15.63 12.57 3.08 17.72
HDMapGen Vector 28.17 $\pm$ 14.81 9.18 13.97 7.02 3.03 2.49 18.10
DiT-L RSI 24.78 $\pm$ 10.38 1.14 7.27 6.11 15.33 1.90 3.95
RLM 32.51 $\pm$ 9.93 1.30 2.36 2.35 3.52 0.88 3.10
DiT-XL* RLM 35.37 $\pm$ 10.28 2.08 3.38 0.27 2.47 0.20 0.47

### v2

@@ Abstract
SLEDGE is the first generative simulator for vehicle motion planning trained on real-world driving logs. Its core component is a learned model that is able to generate agent bounding boxes and lane graphs. The model’s outputs serve as an initial state for rule-based traffic simulation. The unique properties of the entities to be generated for SLEDGE, such as their connectivity and variable count per scene, render the naive application of most modern generative models to this task non-trivial. Therefore, together with a systematic study of existing lane graph representations, we introduce a novel raster-to-vector autoencoder. It encodes agents and the lane graph into distinct channels in a rasterized latent map. This facilitates both lane-conditioned agent generation and combined generation of lanes and agents with a Diffusion Transformer. Using generated entities in SLEDGE enables greater control over the simulation, e.g. upsampling turns or increasing traffic density. Further, SLEDGE can support 500m long routes, a capability not found in existing data-driven simulators like nuPlan. It presents new challenges for planning algorithms, evidenced by failure rates of over 40% for PDM, the winner of the 2023 nuPlan challenge, when tested on hard routes and dense traffic generated by our model. Compared to nuPlan, SLEDGE requires 500$\times$ less storage to set up (<4 GB), making it a more accessible option and helping with democratizing future research in this field.
Keywords:
Diffusion Transformers Simulation Planning Driving
@@ 1 Introduction
While recent breakthroughs in generative AI have revolutionized natural image synthesis [14, 4], generative models are yet to find widespread adoption in autonomous driving. In contrast to the regular pixel grid of images, self-driving planners typically require abstract bird’s eye view (BEV) representations as input which characterize the most important scene elements (e.g., lanes, traffic lights, static and dynamic objects) in a compact, vectorized format (see Fig. 2). These representations are a key component of data-driven simulators which are necessary for rigorous evaluation of planners [10, 49, 24, 9]. However, learning a generative model on such irregular vectorized representations is challenging.
Consequently, many existing data-driven simulators [20, 15] are initialized by simply replaying logs of abstract representations. They extract the local lane graph from a High Definition (HD) map and object bounding boxes from pre-recorded annotation logs. Planning algorithms can be tested in scenarios with restricted routes and durations ($\sim$15 seconds long), to ensure that the environment during simulation is covered in the recording. Moreover, to provide sufficient diversity among routes for comprehensive testing, these simulators require huge databases, e.g., nuPlan [20] consists of 1300 hours of driving logs which require over 2 TB of storage. Such high resource requirements heighten the barrier for entry into the field of vehicle motion planning.
In this paper, we study the generation of simulation-ready abstract representations of driving scenes. Using generative models as an alternative to log replays has the potential for significant compression [38]. However, the unique characteristics of abstract representations in driving scenes pose new challenges to modeling them, e.g., they require accurate topological connectivity, variable entity counts, and precise modeling of geometry (e.g., parallel lines). Due to these characteristics, generative models that operate on uniformly sized representations (e.g., images) are incompatible with our data-driven simulation task.
@@
To tackle these complexities, we first perform a systematic study of the existing representations of lane graphs used in autonomous driving. We then propose a novel representation based on a raster-to-vector autoencoder. It represents a driving scene with a fixed-size BEV rasterized latent map (RLM). We learn to generate these RLMs with a Diffusion Transformer (DiT) [34]. Our model generates high-fidelity results enabling both lane-conditioned agent generation or joint lane and agent generation within a single flexible and scalable framework (Fig. 1). Independent to its use of generative models, SLEDGE also provides the previously missing functionality of simulating only agents within a certain radius of the ego-vehicle. By doing so, we can test on routes that are significantly longer than those currently used for evaluating planners. We find previously ignored failure modes of the state-of-the-art PDM-Closed planner [10], which is unable to complete 25-50% of our new 500m long test routes despite having failure rates below 10% on existing benchmarks.
Contributions. (1) We formalize the task of abstract scene generation for autonomous driving with a challenging benchmark and corresponding metrics. (2) We perform a systematic exploration of modern generative models (with various architectures and representations) and propose a novel latent diffusion model for synthesizing abstract driving scenes that largely outperforms other baselines. (3) We present a simulation framework, SLEDGE, which is nearly 3 orders of magnitude more storage efficient than nuPlan yet enables more rigorous testing of planning algorithms with long-horizon simulations using rule-based traffic.
@@ 2 Related Work
Diffusion Models. Best known for their success in generative modeling of images [37, 14] and video [3, 4, 47], diffusion models have recently found widespread adoption in diverse domains, including point clouds [29, 48, 32], floor plans [40, 6], molecules [46], robot policies [7], traffic patterns [51], and many others. Scenario Diffusion [35], a pioneering approach in generating vehicles conditioned on HD maps with diffusion, is the closest existing approach to ours. This method uses latent diffusion with a raster-based vehicle decoder unlike our proposed transformer decoder head. Importantly, we offer significantly increased capabilities: generation of lane graphs, support for pedestrians, obstacles, and traffic lights, as well as long-horizon simulation environments with reactive agents.
@@
Generating Lane Graphs. Lane graphs, the most important component of HD maps, are well-studied. They are often constructed through an offline mapping process often involving human annotators [13]. However, a surge of recent work on predicting lane graphs from sensor data [25, 23] has sparked interest on generative modeling of these graphs. The first and only existing study on this task, HDMapGen [31], proposes an autoregressive approach for generating lane graphs node-by-node [8]. Our experiments show that this achieves reasonable results, but is unable to match the high quality and scalability offered by our model. Unlike HDMapGen, our approach jointly generates agents with lanes, and efficiently generates all elements in parallel. A concurrent project, DriveSceneGen [41], generates lanes and vehicles with image-space diffusion. Our approach covers agent types beyond vehicles, uses less heuristics, and is more efficient. Additionally, we show the successful integration of our model into a simulator.
@@
Data-driven Simulation. Developing an autonomous driving system necessitates rigorous testing which is costly and risky if conducted in the real world. Driving simulators are an alternative [45, 12, 20]. However, simulators face challenges in ensuring realism while initializing traffic scenes, simulating traffic. Instead, data-driven simulators address these challenges by replaying traffic scenes from real-world recordings [1, 20, 22, 15]. The simulator can mine specific situations or even optimize the initial parameters for safety-critical scenarios [11, 16, 44]. Leveraging modern generative models, we take a step further than existing frameworks and learn the underlying distribution of the real-world data.
3 Method
@@
Our goal is to design a driving scene synthesis framework that can be trained using real-world driving logs and incorporated into SLEDGE, our generative simulator with rule-based traffic. We base this framework on LDMs [37] as: (1) latent diffusion shows excellent training stability and scalability with compute. (2) One can easily construct a fixed-size latent space for diffusion that can be mapped to the variable sized set representation for simulation (Section 3.1) using detection-based transformer architectures [5, 25]. Our LDM is trained in two stages: an autoencoder (Section 3.2) followed by a diffusion model (Section 3.3). We detail the simulation of scenes generated by the LDM in Section 3.4.
@@
We combine sets of entities to represent scenes in the default nuPlan format.
@@ 3.2 Raster-to-Vector Autoencoder
Each entity type in $\mathcal{S}$ is unique. To maintain overall scene consistency, we would like to model them with a single architecture, instead of creating several independent entity-specific generative models. Furthermore, most existing tools in the literature have been developed and optimized for 2D input domains [37]. To this end, we propose the raster-to-vector autoencoder (RVAE) which unifies all entity types in $\mathcal{S}$ into a compact, shared 2D representation well-suited for diffusion modeling. The autoencoder is inspired by work in object detection [5] and online mapping for autonomous driving [25].
@@
Figure 3: Raster-to-Vector Autoencoder (RVAE). We represent scenes with a rasterized latent map (RLM) consisting of two channel groups. The ‘Lanes’ group is decoded into lane segments and the ‘Agents’ group into all other scene entities, via a transformer decoder with attention masking. The autoencoder is trained to predict polylines, bounding boxes, and the ego velocity in a simulation-compatible format.
Vectorization. This step is necessary to decode the unified RSI representation $\mathbf{I}$ back into the per-entity attributes (e.g., polylines, bounding boxes). We use a learned vectorization pipeline consisting of a raster encoder $\pi$ and vector decoder head $\phi$. The ResNet-50 [17] encoder takes the RSI and outputs a rasterized latent map (RLM) $\mathbf{M}=\pi(\mathbf{I})$ of shape $W^{\prime}\times H^{\prime}\times C$, where $H^{\prime}=H/2^{d}$ and $W^{\prime}=W/2^{d}$ for a downsampling factor $d\in\mathbb{N}$. $C$ is a chosen channel dimension. In practice, the RLMs we use are compact, $H^{\prime}=W^{\prime}=8$ and $C=64$. Additionally, as shown in Fig. 3, we split the RLM’s channels into 2 groups, $C=C_{L}+C_{A}$ for the lanes and agents respectively. For each group, we tokenize the latent vectors spatially with $1\times 1$ patches, resulting in $W^{\prime}\times H^{\prime}$ lane tokens and $W^{\prime}\times H^{\prime}$ agent tokens. Following the DETR [5] paradigm, we use these tokens as keys and values for our transformer decoder $\phi$. The decoder uses a fixed number of learnable queries of each entity type, which we cap to a maximum count per entity, based on statistics from our dataset. The final decoder layer is unique per entity type and outputs the attributes specific to that entity, e.g. a $20\times 2$ set of point coordinates for a polyline, or a 6-dimensional descriptor (2D position, orientation, 2D extent, and speed) for a bounding box. It also outputs an existence attribute $p\in[0,1]$ for both polylines ($p_{line}$) and bounding boxes ($p_{box}$), which is used to handle variable counts of ground truth entities with a fixed number of queries.
@@
Training. The autoencoder is optimized using both reconstruction and existence losses, and a KL divergence loss on the RLM. For reconstruction, we first match generated and ground truth entities using the Hungarian algorithm, as in [5]. We use a matching score based on the L1 error of the entity’s position attributes. We then use the L1 error summed over all attributes and averaged over all matches as the training loss. For the existence variable, we use a binary cross entropy loss based on whether the query was matched to a ground truth entity.
@@ 3.3 Diffusion Transformer
Training. For each scene, we sample a noise scaling factor $\sigma$ from a log-normal distribution and create a noisy sample $\hat{\mathbf{M}}=\mathbf{M}+\sigma\boldsymbol{\mathcal{E}}$, with $\boldsymbol{\mathcal{E}}$ sampled from a standard normal distribution of the same shape as $\mathbf{M}$. We model $\delta(\hat{\mathbf{M}};\mathbf{c},\sigma)$ as a Transformer [43], following DiT [34], where $\mathbf{c}$ is a conditioning vector. We choose $\mathbf{c}$ to be a one-hot label indicating the city to which the example belongs. This conditioning resolves ambiguities between locations (e.g. right- and left-hand driving in the US and Singapore). However, it is possible to adapt our framework to other conditioning, e.g., images, text descriptions, or learned clusters based on the autoencoder features. The DiT architecture is simple, scalable, and free from down- or upsampling operations, making it compatible with RLMs of any spatial resolution. It applies a series of self-attention blocks to the tokenized input $\hat{\mathbf{M}}$, with the conditioning on $\mathbf{c}$ and $\sigma$ implemented using AdaLN-Zero [34]. We optimize the L2 reconstruction loss between $\boldsymbol{\mathcal{E}}$ and $\delta(\hat{\mathbf{M}};\mathbf{c},\sigma)$.
@@ 4 Experiments
Implementation. We consider a ResNet-50 for the raster encoder $\pi$ and a transformer decoder with three layers for the vector decoder $\phi$ of the RVAE. For the generative tasks, we apply the DiT-L and DiT-XL (138M vs. 487M params) variants with a $1\times 1$ patch size and DDPM noise scheduling as in [37].
@@ 4.1 Lane Graph Representations
Baselines. We consider three baseline representations. (1) RSI: we use the skan library [33] to extract vector polylines from the $256\times 256\times 2$ RSI representation of the lane graph. Skan is a highly optimized image processing pipeline for graph extraction (details and visualizations in supplementary). (2) RLM w/o mask: we train the RVAE without the channel group masking proposed in Section 3.2, which entangles information about agents and lanes into a single $8\times 8\times 64$ latent map instead of an $8\times 8\times 32$ tensor for the lane graph and an independent $8\times 8\times 32$ tensor for agents. (3) Vector: As an upper bound, we additionally compute our metrics for the respresentation used as target labels for the autoencoder during training. This is a set of polylines shaped $N\times 20\times 2$, where we cap $N$ at 30 in our experiments. Note that the ground truth for evaluation in this experiment uses all polylines in the scene (which is sometimes greater than 30).
@@
Table 2: 64m×$\times$64m Lane Graph Generation. We show the route lengths, precision and recall (in %), and various Frechet distances between samples from the validation set and model outputs. We additionally include qualititve results (more examples in supplementary material). Latent diffusion combining DiT with an RLM obtains the best results. *Trained with $\sim$6$\times$ more compute than others, which already converge at the lower compute budget.
Arch. Repr. Route Prec. (CNN) $\uparrow$ Recall (CNN) $\uparrow$ Frechet (Urban Planning) $\downarrow$
Length $\uparrow$ ImNet RVEnc ImNet RVEnc Conne. Densi. Reach Conve.
VAE RSI 2.68 $\pm$ 3.66 26.58 0.00 8.70 0.16 9.45 0.99 2.86 13.06
RVAE Vector 23.79 $\pm$ 9.96 10.41 4.56 16.28 8.14 15.63 12.57 3.08 17.72
HDMapGen Vector 28.17 $\pm$ 14.81 19.10 7.48 17.17 12.45 7.02 3.03 2.49 18.10
DiT-L RSI 24.78 $\pm$ 10.38 20.36 19.20 21.49 5.94 6.11 15.33 1.90 3.95
RLM 32.51 $\pm$ 9.93 33.25 63.99 36.24 61.60 2.35 3.52 0.88 3.10
DiT-XL* RLM 35.37 $\pm$ 10.28 42.05 78.07 42.32 72.63 0.27 2.47 0.20 0.47